1.      The height of grade 1 pupils is approximately normally distributed with µ = 45 inches and s = 2. a.      If a pupil is selected at random, what is the probability that is taller than 46 inches?


1
Expert's answer
2021-02-28T07:02:30-0500

"Z=\frac{X-\mu}{\sigma}"

"P(X>46)=1-P(X<46)"

"=1-P(Z<\frac{46-45}{2})"

"=1-P(Z<0.5)"

"=1-0.6915=0.3085"